# Autotune: fast, accurate, and automatic tuning parameter selection for LASSO

### v1

@@ 1 Introduction
In high-dimensional regression with the Lasso, model selection is performed by optimally choosing a single tuning parameter $\lambda$, the penalty level. Recent literature has underscored a need for fast and accurate methods to select $\lambda$ [wu2020survey]. This practice of tuning is also in sharp contrast with the classical practice of regression in low-dimensional settings with a handful of predictors, where one first estimates both the regression coefficients ($\beta$) and the noise standard deviation ($\sigma$). Selection of a predictor in the final model is then governed by a tuning parameter $\alpha$, the significance level, which the practitioner can choose based on their perceived cost of false positive in their context. In this work, we adopt an algorithmic view and develop $\mathsf{autotune}$, a strategy for the Lasso to tune itself by relying only on a pre-specified parameter $\alpha\in(0,1)$. The strategy is faster than existing alternatives, and shows better generalization and model selection in our simulation experiments where signal-to-noise ratio is low.
@@
The tuning parameter $\lambda$ controls the level of regularization, and should be chosen to strike a balance between false positives and false negatives in model selection. A faster tuning strategy can help build algorithms that repeatedly uses Lasso, e.g. in the analysis of streaming or sequential data in finance [hallac2017network]. A more efficient tuning strategy that performs better model selection with fewer samples can aid new discoveries in large-scale omics problems in biology [ghaemi2019multiomics].
Among the existing tuning strategies, information criteria such as AIC or BIC are fast, but known to perform poorly in high-dimension [wu2020survey, chen2008extended, zhao2015mixture]. This makes cross-validation (CV) the most popular tuning strategy. But even ignoring its ad hoc features (e.g. randomness in data splitting, arbitrary choice of the number of folds), CV is computationally intensive and not preferable in settings that require multiple Lasso fits, e.g. Gaussian graphical models and Vector Autoregressive (VAR) models. CV is also known to overfit [wang2007tuning, lei2020cross] and experience accuracy loss [bates2024cross] in high-dimensional settings. These issues are more prominent for time series data, where a time series cross-validation (TSCV) is used to select folds while respecting the temporal ordering of samples.
To avoid the challenges of tuning, a number of prior works have proposed changing the loss function of Lasso to make it tuning-free, e.g. Scaled Lasso [sun2012scaled], Square Root Lasso [belloni2011square], Rank Lasso [wang2020tuning]. To date, in high-dimensional problems, none of these methods are as fast as the Lasso implemented in glmnet, and their asymptotic theory for time series data is less developed compared to that of Lasso. For VAR models, [kock2025data] has recently proposed an automatic tuning strategy with theoretical guarantees. In this work, instead of changing the loss function, we ask the following question:
@@
Figure 1: (1(a)) Mean runtimes (with 1se error bars) of $\mathsf{autotune}$, CV and Scaled Lasso as a function of $p$ increasing from 30 to 1500 with $n=300$ and $\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{0}=5$. Fig.˜1(b) Mean runtimes of $\mathsf{autotune}$ VAR, TSCV VAR Lasso, and KPS Lasso ˜23 as a function of number of time series (NN) for VAR(1) estimation with $n=150$. Dimension of regression coefficients $p=N^{2}$, so $p$ here varies from 9 to 900. 1(c) Relative test error (metric on y-axis, for details see ˜34) of solution paths of $\mathsf{autotune}$ (blue, whose endpoint is indicated by the blue star symbol) and Lasso(red) against $\log(\lambda)$ for $n=80,p=750$ and $\mathopen{}\mathclose{{\left\|{\beta_{\mathrm{true}}}}}\right\|_{0}=5$. For Lasso, the best $\lambda$ in it’s path is chosen via CV/AIC/BIC, indicated by the labeled vertical lines. For all experiments, we set $\alpha=0.01$ for $\mathsf{autotune}$.
In Fig.˜1, we highlight the performance of $\mathsf{autotune}$ in light of the above question. Fig.˜1(a) shows that autotune is 15-125 times and 52-190 times faster than CV Lasso and Scaled Lasso respectively over a number of simulation scenarios. The speed gains of $\mathsf{autotune}$ become more pronounced in VAR Lasso, where multiple Lasso’s need to be run and tuning is done via TSCV. Fig.˜1(b) shows that $\mathsf{autotune}$ VAR is 73-186 times faster than VAR Lasso tuned via TSCV and 43-97 times faster than the VAR Lasso of [kock2025data] with automatic tuning. In Fig.˜1(c), we illustrate the reason of its speed on a simulated data with $n=80,p=750$, and $\mathopen{}\mathclose{{\left\|{\beta_{\mathrm{true}}}}}\right\|_{0}=5$. Instead of evaluating Lasso solutions on a dense grid of $\lambda$ (red dots), $\mathsf{autotune}$ (blue dots/line) searches over only 3 different data-adaptively chosen $\lambda$ values before stopping (blue star). Interestingly, $\mathsf{autotune}$ picks a $\lambda$ that is nearly optimal for relative test error (RTE, see 34), while AIC, BIC overfit and CV min ˜($\lambda$1) and 1se ˜($\lambda$2) rule both underfit.
@@
Our $\mathsf{autotune}$ alternately estimates $\beta$ and $\sigma$ in a full negative Gaussian log-likelihood with an $\ell_{1}$ penalty:
Since for any given $\sigma^{2}$, solving for $\beta$ in (3) is essentially Lasso with tuning parameter $\lambda:=\lambda_{0}\sigma^{2}$, sequentially guessing different values of $\sigma^{2}$ results in fitting Lasso on an automatic, data-driven sequence of $\lambda$ values.
These estimators of $\sigma$ are expected to suffer from the shrinkage bias of $\widehat{\beta}$ throughout the iteration. Our main technical contribution is to use the partial residuals (PR) to correct such shrinkage bias of $\widehat{\sigma}^{2}$ efficiently throughout the iterations.
For a given guess of $\sigma$, and hence a given $\lambda$, the coordinate descent (CD) algorithm for Lasso updates the regression coefficients iteratively using the formulae $\widehat{\beta}_{k}=S_{\lambda}(\sum_{i=1}^{n}\text{x}_{ij}r^{(k)}_{i}/n)$, for $k=1,2,\ldots,p$, where $S_{\lambda}(.)$ is the soft thresholding operator and the coordinates of the partial residuals
$r^{(k)}$ of the $k^{th}$ predictor are defined as
[…]
We observe that for sparse $\beta$, the vector of $p$ standard deviations of PR, $SD(r^{(k)})$, $k=1,2,\ldots,p$, exhibit an interesting behavior throughout the CD iterations. Specifically, if $\beta$ has $s$ non-zero entries, about $s$ many coordinates in the above vector are outliers, i.e. unusually larger than the rest. So, we build a linear model of Y by sequentially introducing these outlier predictors one at a time, using a classical F-test at a significance level $\alpha$ to determine when to stop. The mean squared error (MSE) of this final linear model is used as the new estimator of $\sigma^{2}$. Since we don’t need the multiple linear regression coefficients $\beta$ themselves and only the $\sigma^{2}$, the computation reduces to one simple Gram-Schmidt orthogonalization, which is much faster than fitting $s$ many linear models. Also, in most cases, this estimator $\sigma^{2}$ converges way before the full CD for $\beta$ converges. So, with this new guess of $\sigma^{2}$, autotune moves on to try a new $\lambda$ before waiting for CD to converge in the current $\lambda$. Only when the sequence of $\widehat{\sigma}^{2}$ converges, leading to a final $\lambda$, we compute our final Lasso solution using CD. This offers substantial gain in speed over methods that require first fitting Lasso over a grid of $\lambda$ values.
@@
In a wide range of simulation experiments using data generating process from [hastie2020best], we benchmark $\mathsf{autotune}$ against CV-tuned and Scaled Lasso. In Sec.˜4.1.2, $\mathsf{autotune}$ shows higher accuracy than both CV-tuned Lasso and Scaled Lasso in terms of estimation (Relative MSE), prediction (“Relative Test Error” [Sec. 3.1, hastie2020best]), and variable selection (Area under ROC curve). In Sec.˜4.1.3, $\mathsf{autotune}$ demonstrates 15-200 times speedup over CV-tuned Lasso and Scaled Lasso. In our simulations for VAR models in Sec.˜4.2, we find that accuracy gains of $\mathsf{autotune}$ over benchmarks on the same metrics are even larger, and $\mathsf{autotune}$ is even faster over TSCV-tuned VAR Lasso and [kock2025data]. Lastly, we show that $\mathsf{autotune}$ improves $\sigma^{2}$ estimation in high-dimensional data as compared to the state-of-the-art methods, leading to better high dimensional inference [dezeure2015high] in Sec.˜4.3.1 and provides promising diagnostics for checking the sparsity assumption in Sec.˜4.3.2.
We also perform experiments on a real-life dataset sp500 (given in Scaled Lasso’s R package, scalreg [scalreg]) in Sec.˜5, and demonstrate the gains of $\mathsf{autotune}$ Lasso over the benchmarks (including Scaled Lasso) in terms of relative test error, Out-of-sample R2 and variable selection. The other sections of the paper are organized as follows. Sec.˜2 explains the prior state-of-the-art algorithms for computing Lasso and tuning $\lambda$(most of which are used as benchmarks for measuring $\mathsf{autotune}$’s performance). Sec.˜3 elucidates the algorithm of $\mathsf{autotune}$ Lasso, $\mathsf{active}$ $\mathsf{autotune}$ Lasso, and $\mathsf{autotune}$ VAR.
@@ 2 Background
In this section, we describe existing tuning strategies for Lasso in regression and VAR modeling, and discuss their limitations. Then we briefly describe partial residuals (PR) in CD algorithms for Lasso, and sequential F-tests for model selection in classical statistics - two key technical ingredients used in $\mathsf{autotune}$.
@@
Tuning $\lambda$ via K-fold cross-validation [hastie2015statistical] remains the most common tuning strategy among practitioners. Denote $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)=\mathop{\mathrm{argmin}}_{\beta\in\real^{p}}\mathcal{L}_{\lambda}(\beta)$, (see ˜2). Once we have $\widehat{\beta}(\lambda)$ on a grid of $\lambda$ values, there are two choices
@@
[…]
@@
Despite its popularity, the number of folds is set to 5 or 10 in an ad hoc manner. The final chosen $\lambda$ ˜($\lambda$1) is also known to be susceptible to the fold assignment used in CV [bovelstad2007predicting, roberts2014stabilizing]. Recent works have also shown that CV overfits the data [wang2007tuning, lei2020cross] and loses accuracy [bates2024cross]. To the best of our knowledge, there has been no systematic study to assess the quality of CV(1se) ˜($\lambda$2) solutions.
@@
AIC [akaike2003new] and BIC [schwarz1978estimating] introduced the idea of penalizing complexity of the model fitted for inducing model selection. The general formula is $n\log\mathopen{}\mathclose{{\left\|{\text{Y}-\text{X}\widehat{\beta}}}}\right\|_{2}^{2}+kp$ where $k=2$ for AIC and $k=\log(n)$ for BIC. For tuning $\lambda$ in Lasso, AIC and BIC’s simplicity, ease of computation, and avoiding repeated data-splitting calculations made them a formidable and popular alternative to CV. But these information criteria were formulated under the assumption that $p$ is fixed and sample size $n\to\infty$, making them unsuitable for high dimensional regime $n<p$ [chen2008extended, zhao2015mixture]. Other information criterion like eBIC [chen2008extended], HBIC [zhao2015mixture] are also proposed to combat scenarios with $p$ growing with $n$ but are not used widely by practitioners and recent research [10.1093/jjfinec/nbab023] suggests they don’t improve upon the performance of BIC in time series analysis.
Linking $\lambda$ with error variance
High-dimensional asymptotics of Lasso for linear model with Gaussian errors suggests choosing $\lambda$ in the order of $\sigma\sqrt{\log(p)/n}$ [bickel2009simultaneous]. Since $\sigma$ is unknown in practice, a line of work reformulate the loss function ˜2 or ˜3 to make the tuning of $\lambda$ independent of $\sigma$ [belloni2011square, lederer2015don, wang2020tuning].
Another line of work, related more to our method, focuses on estimating $\sigma$ from data and using it for tuning $\lambda$. For example, Scaled Lasso [sun2012scaled] introduced a $\sigma$-based tuning procedure for $\lambda$ by modifying the penalized Gaussian log-likelihood objective ˜3.
@@
˜5 was motivated by the discussions [sun2010comments, antoniadis2010comments] on the paper [stadler2010]. Scaled Lasso [sun2012scaled] proposes an iterative minimization algorithm alternating over $\beta$ and $\sigma^{2}$ for tuning $\lambda$ in Lasso: use the $\widehat{\beta}$’s from the Lasso solution path to calculate the RSS for estimating the $\sigma$, and then the $\widehat{\sigma}$ tunes the $\lambda$ for the next iteration of coordinate descent. Moreover, [sun2012scaled] derived asymptotic guarantees of Scaled Lasso for estimating $\sigma$.
Additionally, [yu2019estimating] proposed Natural and Organic Lasso which optimized the following respective modified loss functions
@@
[yu2019estimating] arrived at Natural Lasso by making the original Lasso objective ˜2 scale invariant. Natural Lasso still needs CV to tune its $\lambda$.
But it’s known that $\widehat{\beta}$ from Lasso’s coordinate descent has shrinkage bias [hastie2015statistical] and using its corresponding RSS will lead to $\widehat{\sigma}$ with large finite-sample errors. Our simulations(Sec.˜4.3) report the same finite sample difficulties after running an extensive simulation study on $\sigma$ estimation in Lasso regression. Finite sample limitations of Scaled Lasso was also reported in [yu2019estimating]’s simulations.
@@
Vector Autoregression (VAR) is a workhorse model in multivariate time series data, widely used in macroeconomics [banbura2010large, bernanke2005measuring], portfolio selection in finance [fan2011sparse], identifying gene regulatory networks in genomics [michailidis2013autoregressive, broll2025prolong] and for understanding functional connectivity between different brain regions in neuroscience [smith2012future, seth2013granger]. Formally, a VAR($d$) model on a $p$-dimensional stationary time series $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}=\mathopen{}\mathclose{{\left\{\mathopen{}\mathclose{{\left(Z^{t}_{1},\dots,Z^{t}_{p}}}\right)}}\right\}_{t\geq 1}$ with no contemporaneous association takes the form
$\mathrm{VAR}\mathopen{}\mathclose{{\left(d}}\right):Z^{t}=\sum_{i=1}^{d}A_{i}Z^{t-i}+\epsilon^{t},\quad\epsilon^{t}\stackrel{{\scriptstyle\textrm{i.i.d.}}}{{\sim}}\mathopen{}\mathclose{{\left(0,\Sigma_{\epsilon}}}\right),\quad\Sigma_{\epsilon}=\mathop{\mathrm{diag}}\mathopen{}\mathclose{{\left(\sigma^{2}_{1},\dots,\sigma^{2}_{p}}}\right),$ (7)
where $A_{1},\dots,A_{d}$ are $p\times p$ transition matrices which need to be estimated. Given $n+d$ stationary observations of $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$, all the $d$ VAR transition matrices can be estimated in the autoregressive design [hyndman2018forecasting]
@@
where $\Phi$ is the $dp\times p$ matrix constructed by stacking $A_{1}^{\top},A_{2}^{\top},\dots,A_{d}^{\top}$. The parameter space dimension ($dp^{2}$) grows quadratically with the number of time series $p$, and the OLS estimate of $\Phi$ will be consistent only if $n>dp^{2}$, which is often a big ask for sample size in modern VAR applications. So, it is common to assume $\Phi$ is sparse and use $\ell_{1}$-penalized least squares regressions [basu2015regularized]. We obtain the VAR Lasso estimates of $\Phi$ by conducting $p$ separate, columnwise Lasso regressions. Mathematically,
@@
A common approach to tune these $p$ different $\lambda_{i}$’s is the time series CV (TSCV) [Sec. 5.10, hyndman2018forecasting]. It is more constrained because the temporal order of the data must be preserved when choosing the different folds. A “rolling window” is adopted, where the dataset is divided into $2K$ folds along the temporal dimension. Then the $(K+1)^{th}$ fold is assigned the role of a test set, and all the folds corresponding to previous time points are used as the training data to fit the VAR Lasso on the original grid of $\lambda_{i}$s. We repeat this process $K$ times, where every fold in the latter half plays the role of test fold once, and in each repetition, all the previous folds are used as training data. Lastly, TSCV averages the $K$ mean squared errors (MSE) for each $\lambda_{i}$ to get the mean TSCV error, and chooses a $\lambda_{i}$ which minimizes this. As a result, TSCV has some statistical limitations over the usual CV [bergmeir2012use, bergmeir2018note]. For example, constraints in data-splitting for TSCV reduces the effective sample size available for tuning of $\lambda$.
Moreover, running separate TSCV on each column can be computationally prohibitive for analysis of modern day large time series datasets. Hence, prior software [bigtimeRpackage] have often focused on uniform penalization across all the separate column-wise lasso regressions, which is inadequate for handling heterogeneity across multiple time series.

### v2

@@ 1 Introduction
In high-dimensional regression with the Lasso, model selection is performed by optimally choosing a single tuning parameter $\lambda$, the penalty level which is expected to strike a balance between false positive and false negatives in model selection. Recent works have underscored a need for fast and accurate methods to select $\lambda$ [wu2020survey]. This practice of tuning is also in sharp contrast with the classical practice of regression in low-dimensional settings with a handful of predictors, where one first estimates both the regression coefficients ($\beta$) and the noise standard deviation ($\sigma$). Selection of a predictor in the final model is then governed by a tuning parameter $\alpha$, the significance level, which the practitioner can choose based on their perceived cost of false positive in their context. In this work, we adopt an algorithmic view and develop $\mathsf{autotune}$, a strategy for the Lasso to tune itself by relying only on a pre-specified parameter $\alpha\in(0,1)$ and jointly estimating $\beta$ and $\sigma$. The strategy is faster than existing alternatives, and shows better generalization and model selection in our simulation experiments where signal-to-noise ratio is low.
@@
The tuning parameter $\lambda$ controls the level of regularization. A faster tuning strategy can help build algorithms that repeatedly uses Lasso, e.g. in the analysis of streaming or sequential data in finance [hallac2017network]. A more efficient tuning strategy that performs better model selection with fewer samples can aid new discoveries in large-scale omics problems in biology [ghaemi2019multiomics].
Among the existing tuning strategies, information criteria such as Akaike and Bayesian Information Criterion (AIC [akaike2003new] and BIC [schwarz1978estimating]) are fast, but known to perform poorly in high-dimension [chen2008extended, zhao2015mixture, wu2020survey]. This makes cross-validation (CV) the most popular tuning strategy. But even ignoring its ad hoc features (e.g. randomness in data splitting, arbitrary choice of the number of folds), CV is computationally intensive and not preferable in settings that require multiple Lasso fits, e.g. Gaussian graphical models and Vector Autoregressive (VAR) models. CV is also known to overfit [wang2007tuning, lei2020cross] and experience accuracy loss [bates2024cross] in high-dimensional settings. These issues are more prominent for time series data, where a time series cross-validation (TSCV) is used to select folds such that the training folds can only contain the observations that occurred prior to the test data points.
To avoid the challenges of tuning, a number of prior works have proposed changing the loss function of Lasso to make it tuning-free, e.g. Scaled Lasso [sun2012scaled], Square Root Lasso [belloni2011square], Rank Lasso [wang2020tuning]. To date, in high-dimensional problems, none of these methods are as fast as the Lasso implemented in glmnet [glmnet], and their asymptotic theory for time series regressions is less developed compared to that of Lasso. For VAR models, [kock2025data] has recently proposed an automatic tuning strategy with theoretical guarantees. In this work, instead of changing the loss function, we ask the following question:
@@
Figure 1: (1(a)) Mean runtimes (with 1 standard error bars) of $\mathsf{autotune}$, CV Lasso and Scaled Lasso as a function of $p$ increasing from 30 to 1500 ($n=300$, $\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{0}=5$). Fig.˜1(b) Mean runtimes of VAR Lasso tuned with $\mathsf{autotune}$, TSCV and KPS Lasso ˜23 against the number of time series (NN) for a VAR(1) model with $n=150$. Dimension of regression coefficients $p=N^{2}$, so $p$ here varies from 9 to 900. (1(c)) Relative test error (metric on y-axis, for details see ˜33) of solution paths of $\mathsf{autotune}$ (blue) and Lasso (red) against $\log(\lambda)$ ($n=80,p=750$ and $\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{0}=5$). We set $\alpha=0.01$ in $\mathsf{autotune}$ for all experiments.
In Fig.˜1, we highlight the performance of $\mathsf{autotune}$ in light of the above question. Fig.˜1(a) shows that for regression, $\mathsf{autotune}$ is 15-125 times and 52-190 times faster than CV Lasso and Scaled Lasso respectively over a number of simulation scenarios. The speed gains of $\mathsf{autotune}$ become more pronounced in high-dimensional VAR, which needs multiple Lasso fits and tuning is primarily done via TSCV. Fig.˜1(b) shows that $\mathsf{autotune}$ VAR is 75-190 times faster than VAR Lasso tuned via TSCV and 45-100 times faster than [kock2025data]. In Fig.˜1(c), we illustrate the reason of its speed on a simulated data with $n=80,p=750$, and $\mathopen{}\mathclose{{\left\|{\beta}}}\right\|_{0}=5$. Instead of evaluating Lasso solutions on a dense grid of $\lambda$ (red dots), $\mathsf{autotune}$ (blue dots) searches over only 3 different data-adaptively chosen $\lambda$ values before stopping (blue star). Interestingly, $\mathsf{autotune}$ picks a $\lambda$ that is nearly optimal for relative test error (RTE, see (33) or [Sec. 3.1, hastie2020best]), while AIC, BIC overfit and the two different tuning rules of CV Lasso underfit.
[…]
We repeated the experiment of Fig.˜1(c) 2500 times to convey that the results in Fig.˜1(c) are not anecdotal. We compare the quality of tuning of $\mathsf{autotune}$ and CV Lasso in Fig.˜2 across 2500 replications, other tuning strategies were not competitive. The $\lambda$ values selected by $\mathsf{autotune}$ are more concentrated around lower RTE region of the $\lambda$ grid than the CV Lasso ˜($\lambda$1). A paired t-test between their RTE across each replication shows $\mathsf{autotune}$ has significantly smaller RTE than CV Lasso (p-value $1.06\times 10^{-14})$.
Essentially, $\mathsf{autotune}$ alternately estimates $\beta$ and $\sigma$ in a full negative Gaussian log-likelihood with an $\ell_{1}$ penalty:
Since for any given $\sigma^{2}$, solving for $\beta$ in (3) is essentially solving Lasso with tuning parameter $\lambda:=\lambda_{0}\sigma^{2}$, sequentially guessing different values of $\sigma^{2}$ results in fitting Lasso on an automatic, data-driven sequence of $\lambda$ values.
These estimators of $\sigma$ are expected to suffer from the shrinkage bias of $\widehat{\beta}$ throughout the iterations. Our main technical contribution is to leverage the partial residuals (PR) to correct such shrinkage bias of $\widehat{\sigma}^{2}$ efficiently in the process of solving the Lasso problem.
For a given guess of $\sigma$, and hence a given $\lambda$, the coordinate descent (CD) algorithm for Lasso updates the regression coefficients iteratively using the formulae $\widehat{\beta}_{k}=\mathcal{S}_{\lambda}(\sum_{i=1}^{n}\text{x}_{ij}r^{(k)}_{i}/n)$, for $k=1,2,\ldots,p$, where $\mathcal{S}_{\lambda}(a)=\mathrm{sign}(a)(|a|-\lambda)_{+}$ is the soft thresholding operator and the coordinates of PR $r^{(k)}$ of the $k^{th}$ predictor are defined as
@@
Since we don’t need the multiple linear regression coefficients $\beta$ themselves and only the $\sigma^{2}$, the computation reduces to one simple Gram-Schmidt orthogonalization, which is much faster than fitting $s$ many linear models. Also, in most cases, this estimator $\sigma^{2}$ converges way before the full CD for $\beta$ converges. So, with this new guess of $\sigma^{2}$, $\mathsf{autotune}$ stops its CD early and goes to a new $\lambda$ rather than waiting for its CD to converge in the current $\lambda$. Only when the sequence of $\widehat{\sigma}^{2}$ converges, leading to a final $\lambda$, we compute our final Lasso solution using CD. This offers substantial gain in speed over methods that require first fitting Lasso over a grid of $\lambda$ values.
@@
In a wide range of simulation experiments using data generating process from [hastie2020best], we benchmark $\mathsf{autotune}$ against CV Lasso and Scaled Lasso. In Sec.˜4.1.1, $\mathsf{autotune}$ shows higher accuracy than both CV Lasso and Scaled Lasso in terms of estimation (Relative Mean Squared Error), prediction (RTE), and variable selection (Area under Receiver Operating Characteristic curve). In Sec.˜4.1.2, $\mathsf{autotune}$ demonstrates 15-200 times speedup over CV Lasso and Scaled Lasso. In our simulations for VAR models in Sec.˜4.2, we find that accuracy gains of $\mathsf{autotune}$ over benchmarks on the same metrics are even larger, and $\mathsf{autotune}$ is even faster over TSCV and [kock2025data]. Lastly, we show that $\mathsf{autotune}$ improves $\sigma^{2}$ estimation in high-dimensional data as compared to the state-of-the-art methods, leading to better high dimensional inference [dezeure2015high] in Sec.˜4.3.1 and provides promising diagnostics for checking the sparsity assumption in Sec.˜4.3.2. In Sec.˜5, we perform experiments on a real-life dataset sp500 (given in the Scaled Lasso R package scalreg [scalreg] and used in [Sec. 6, wu2020survey]), and $\mathsf{autotune}$ generalizes better in test data with sparser model as compared to CV Lasso and Scaled Lasso.
The rest of this paper is organized as follows. Sec.˜2 explains the prior state of the art algorithms for computing Lasso and tuning $\lambda$, and describes some technical ingredients used in $\mathsf{autotune}$. Sec.˜3 describes the algorithms of $\mathsf{autotune}$ Lasso for regression and VAR. Sec.˜4 report the performance of $\mathsf{autotune}$ on simulated datasets against benchmarks described in Sec.˜2. Lastly, we compare the performance of $\mathsf{autotune}$ against the becnhmarks on a real-life financial dataset in Sec.˜5.
@@ 2 Background
In this section, we describe existing tuning strategies for Lasso in regression and VAR modeling, and discuss their limitations. Then we briefly describe PR in CD algorithms for Lasso, and sequential F-tests for model selection in classical statistics - two key technical ingredients used in $\mathsf{autotune}$.
@@
Tuning $\lambda$ via K-fold cross-validation [hastie2015statistical] remains the most common tuning strategy among practitioners. Denote $\widehat{\beta}\mathopen{}\mathclose{{\left(\lambda}}\right)=\mathop{\mathrm{argmin}}_{\beta\in\real^{p}}\mathcal{L}_{\lambda}(\beta)$, (see ˜2). Once we have $\widehat{\beta}(\lambda)$ on a grid of $\lambda$ values, there are two choices in $\mathsf{glmnet}$ [glmnet]:
@@
    CV(min): $\lambda$ which minimized the K-fold CV error among all the $\lambda$ values in the $\lambda$-grid,
@@
Despite its popularity, the number of folds in CV is set to 5 or 10 in an ad hoc manner. The final chosen $\lambda$ ˜($\lambda$1) is also known to be susceptible to the fold assignment used in CV [bovelstad2007predicting, roberts2014stabilizing]. Recent works have also shown that CV overfits the data [wang2007tuning, lei2020cross] and loses accuracy [bates2024cross]. To the best of our knowledge, there has been no systematic study to assess the quality of CV(1se) ˜($\lambda$2) solutions.
@@
AIC [akaike2003new] and BIC [schwarz1978estimating] introduced the idea of penalizing complexity of the model fitted for inducing model selection. The general formula is $n\log\mathopen{}\mathclose{{\left\|{\text{Y}-\text{X}\widehat{\beta}}}}\right\|_{2}^{2}+kp$ where $k=2$ for AIC and $k=\log(n)$ for BIC. For tuning $\lambda$ in Lasso, their ease of computation, and avoiding repeated data-splitting calculations made them a simpler alternative to CV. But these information criteria were formulated under the assumption that $p$ is fixed and sample size $n\to\infty$, making them unsuitable for high-dimensional regime $n<p$ [chen2008extended, zhao2015mixture]. Other information criterion like eBIC [chen2008extended], HBIC [zhao2015mixture] are also proposed for scenarios where $p$ grows with $n$. But these are not yet adopted widely and some recent research [10.1093/jjfinec/nbab023] suggests that they do not improve upon the performance of BIC in time series analysis.
Linking $\lambda$ with error variance:
High-dimensional asymptotics of Lasso for linear model with Gaussian errors suggests choosing $\lambda$ in the order of $\sigma\sqrt{\log(p)/n}$ [bickel2009simultaneous]. Since $\sigma$ is unknown in practice, a line of work reformulates the loss function ˜2 or ˜3 to make the tuning of $\lambda$ independent of $\sigma$ [belloni2011square, lederer2015don, wang2020tuning]. Statistical properties of these methods in high-dimensional time series are still topics ongoing research. In this work, we instead focus on tuning Lasso whose statistical properties are by now well-understood for time series data.
Another line of work, related more to our method, focuses on first estimating $\sigma$ from data and then using it for tuning $\lambda$. For example, Scaled Lasso [sun2012scaled] introduced a $\sigma$-based tuning procedure for $\lambda$ by modifying the penalized Gaussian log-likelihood objective ˜3.
@@
In fact, ˜5 was motivated by the discussions [sun2010comments, antoniadis2010comments] on the paper [stadler2010]. Scaled Lasso [sun2012scaled] proposes an iterative minimization algorithm alternating over $\beta$ and $\sigma^{2}$ for tuning $\lambda$ in Lasso: use the $\widehat{\beta}$ values from the Lasso solution path to calculate the Mean Squared Error (MSE), then do a degrees of freedom correction on MSE to estimate the $\sigma$, and then the $\widehat{\sigma}$ tunes the $\lambda$ for the next iteration in Scaled Lasso. Moreover, [sun2012scaled] derived asymptotic guarantees of $\widehat{\sigma}$ of Scaled Lasso.
Additionally, [yu2019estimating] proposed Natural and Organic Lasso which optimized the following modified loss functions respectively:
@@
[yu2019estimating] arrived at Natural Lasso by making the original Lasso objective ˜2 scale invariant. Natural Lasso still needs CV to tune its $\lambda$. For Organic Lasso, authors [Thm. 12, yu2019estimating] gave a theoretical choice of $\lambda$ using the asymptotic theory of Lasso.
But it is known that $\widehat{\beta}$ from Lasso has shrinkage bias [hastie2015statistical], and using its corresponding MSE may introduce that bias in estimating $\widehat{\sigma}$, leading to inaccurate tuning. While such bias correction is usually done in the literature after the final Lasso fit using ordinary least squares [MEINSHAUSEN2007374, postlasso], an attempt to correct or even reduce such shrinkage bias in $\widehat{\sigma}$ during the Lasso fit may offer some advantage in tuning. We attempt to achieve this in $\mathsf{autotune}$.
@@
Vector autoregression (VAR) is a workhorse model in multivariate time series data, widely used in macroeconomics [banbura2010large, bernanke2005measuring], portfolio selection in finance [fan2011sparse], identifying gene regulatory networks in genomics [michailidis2013autoregressive, broll2025prolong] and for understanding functional connectivity between different brain regions in neuroscience [smith2012future, seth2013granger]. Formally, a VAR($d$) model on a $p$-dimensional stationary time series $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$ with no contemporaneous association takes the form
$\mathrm{VAR}\mathopen{}\mathclose{{\left(d}}\right):Z^{t}=\sum_{i=1}^{d}A_{i}Z^{t-i}+\varepsilon^{t},\quad\varepsilon^{t}\stackrel{{\scriptstyle\textrm{i.i.d.}}}{{\sim}}\mathopen{}\mathclose{{\left(0,\Sigma_{\varepsilon}}}\right),\quad\Sigma_{\varepsilon}=\mathop{\mathrm{diag}}\mathopen{}\mathclose{{\left(\sigma^{2}_{1},\dots,\sigma^{2}_{p}}}\right),$ (7)
where $A_{1},\dots,A_{d}$ are $p\times p$ transition matrices which need to be estimated and $\sigma^{2}_{1},\dots,\sigma^{2}_{p}$ are the noise variances. Given $n+d$ stationary observations of $\mathopen{}\mathclose{{\left\{Z^{t}}}\right\}_{t\geq 1}$, all the $d$ VAR transition matrices can be estimated in the autoregressive design [hyndman2018forecasting]
@@
where $\Phi$ is the $dp\times p$ matrix constructed by stacking $A_{1}^{\top},A_{2}^{\top},\dots,A_{d}^{\top}$. The parameter space dimension ($dp^{2}$) grows quadratically with the number of time series $p$, and the OLS estimate of $\Phi$ will be consistent only if $n>dp^{2}$, which is uncommon in modern VAR applications. So, it is common to assume $\Phi$ is sparse and use $\ell_{1}$-penalized least squares regressions [basu2015regularized]. We obtain the VAR Lasso estimator of $\Phi$ by conducting $p$ separate, columnwise Lasso regressions. Mathematically,
@@
A common approach to tune these $p$ different values of $\lambda_{i}$ is TSCV [Sec. 5.10, hyndman2018forecasting]. It is more constrained because the temporal order of the data must be preserved when choosing the different folds. A “rolling window” is adopted, where the dataset is divided into $2K$ folds along the temporal dimension. Then the $(K+1)^{th}$ fold is assigned the role of a test set, and all the folds corresponding to previous time points are used as the training data to fit the VAR Lasso on the original grid of $\lambda_{i}$ separately for each $i\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$. We repeat this process $K$ times, where every fold in the latter half plays the role of test fold once, and in each repetition, all the previous folds are used as training data. Lastly for each $i\in\mathopen{}\mathclose{{\left\{1,\dots,p}}\right\}$, TSCV averages the $K$ mean squared errors (MSE) for every value in the grid of $\lambda_{i}$ to get the mean TSCV error, and chooses a $\lambda_{i}$ which minimizes this. As a result, TSCV has some statistical limitations over the usual CV [bergmeir2012use, bergmeir2018note]. For example, constraints in data-splitting for TSCV reduces the effective sample size available for tuning of $\lambda$.
Moreover, running $p$ separate TSCV on each column can be computationally prohibitive for analysis of modern day large time series datasets. Hence, prior software [bigtimeRpackage] have often focused on uniform penalization across all the separate column-wise lasso regressions, which is inadequate for handling heterogeneity across multiple time series.